# LESS: Label-Efficient and Single-Stage Referring 3D Segmentation

### v1

@@ Abstract
Referring 3D Segmentation is a visual-language task that segments all points of the specified object from a 3D point cloud described by a sentence of query. Previous works perform a two-stage paradigm, first conducting language-agnostic instance segmentation then matching with given text query. However, the semantic concepts from text query and visual cues are separately interacted during the training, and both instance and semantic labels for each object are required, which is time consuming and human-labor intensive. To mitigate these issues, we propose a novel Referring 3D Segmentation pipeline, Label-Efficient and Single-Stage, dubbed LESS, which is only under the supervision of efficient binary mask. Specifically, we design a Point-Word Cross-Modal Alignment module for aligning the fine-grained features of points and textual embedding. Query Mask Predictor module and Query-Sentence Alignment module are introduced for coarse-grained alignment between masks and query. Furthermore, we propose an area regularization loss, which coarsely reduces irrelevant background predictions on a large scale. Besides, a point-to-point contrastive loss is proposed concentrating on distinguishing points with subtly similar features. Through extensive experiments, we achieve state-of-the-art performance on ScanRefer dataset by surpassing the previous methods about 3.7% mIoU using only binary labels.
@@ 1 Introduction
Previous Referring 3D Segmentation [13; 27] mainly leverage a two-stage workaround, as shown in Fig.1 (a). They typically adopt a 3D instance semantic segmentation network to get the instance proposals at the first stage. The predicted instance proposals will be utilized to match with the queries and finally get the final prediction mask according to the matching score. Although those method have achieved remarkable performance, there still exist some problems. First, owing to the large-scale and irregular 3D point clouds, some instance proposals may leave out the target in the pre-segmentation stage. Besides, lacking of linguistic guidance in the segmentation stage fails to focus on the objects that are more essential to the referring task. Moreover, existing Referring 3D Segmentation utilizes both instance labels and semantic labels to segment target proposal rather than binary mask used in referring image segmentation, which is more time consuming and labor intensive.
@@ 2.1 Referring 3D Segmentation
3D visual grounding aims to locate the object within point clouds mentioned by free-form natural language descriptions. Most methods follow a two-stage detection-then-matching pipeline. Initially, they utilize a pre-trained 3D detector [26; 21] or segmenter [15; 31] to extract object representations. Subsequently, these methods align text features with object features to identify the best-matched object. Researchers primarily concentrate on the second stage which involves modeling object relationships and exploring feature fusion between language and objects. Methods employed include multi-modal feature concatenation [3; 1], attention-based multi-modal feature alignment [41; 11], graph neural network-based reasoning [13; 38], and the alignment of visual and language features aided by 2D images [37; 39; 35].Other researchers have investigated single-stage approaches for 3D visual grounding. 3D-SPS [23] views the task as key point selection, progressively identifying keypoints with the guidance of language and directly locates the target. BUTD-DETR [14] employs a transformer decoder [2] to identify described objects using language cues and proposal boxes. Building on this, EDA [34] enhances dense alignment between objects and point clouds by explicitly decoupling textual attributes from sentences.
@@ 2.3 Referring Image Segmentation
Referring image segmentation is a visual task that involves pixel-level segmentation of an image’s target object based on a referring expression. Early approaches [12; 18; 4] utilize CNNs and RNNs to extract features and then perform simple concatenation for multi-modal feature fusion. Subsequent research [16; 17; 9] adopt transformer models for more effective feature extraction and fusion. Recent studies [36; 10; 40] focus on identifying optimal positions for language-vision alignment. Additionally, some methods [32; 5] have enhanced alignment between language and pixel features by designing specialized loss functions. Other approaches [19; 42] treat referring image segmentation as an auto-regressive vector generation problem, creating masks from generated closed vectors.
@@ 3.1 Visual and Text Feature Extractor
The point cloud scene $P\in\mathbb{R}^{N\times 6}$ contains $N$ points in the scene, and each point is represented with six dimensions of RGBXYZ attributes. We first voxelize points into regular voxels and adopt a sparse 3D U-Net [29; 15] to extract point-wise fused feature $F\in\mathbb{R}^{N\times C}$. Here the encoder part of the U-Net has $5$ stages and the feature from the $i$-th encoder stage is denoted as $V_{i}$.
@@
Due to the lack of such rich annotations as [13; 27], it is crucial for our model to learn the relationship between fine-grained word-level features and point-level features in such a point-level segmentation task. Meanwhile, we notice that leveraging the rich convolutional layers in the encoder to excavate multi-modal context is effective way to extract language-aware visual feature. Therefore, we design Point-Word Cross-Modal Alignment (PWCA) module, as shown in the lower left part in Fig.2, which contains a standard cross-attention module [30] and an MLP-Tanh gate. Cross-attention module aligns point-wise and word-wise feature to get the language-aware visual feature. A nonlinear tanh gate is adopted to prevent the fused signal from overwhelming the original signal. Given the multi-scale point cloud features $\{V_{i}\in\mathbb{R}^{N_{i}\times C_{i}}\}_{i=1}^{5}$, we simply project the word feature $W\in\mathbb{R}^{L\times C}$ to $W_{i}\in\mathbb{R}^{L\times C_{i}}$. PWCA can be formulated as follow:
@@ 3.2 Query Mask Predictor
Inspired by [29; 28], Query Mask Predictor (QMP) module, as shown in Fig.2, takes fused feature $F\in\mathbb{R}^{N\times C}$ and learnable queries $Q_{0}\in\mathbb{R}^{K\times C}$ as input and progressively distinguishes the referring target by multi-layer cross-modal transformers. Finally, we extract the proposal masks $M_{m}\in\mathbb{R}^{K\times N}$ based on queries embeddings $Q_{m}\in\mathbb{R}^{K\times C}$ and fused feature $F$.
@@ 3.4 Loss Function
Different from previous work [13; 27], we take the Referring 3D Segmentation task as segmentation task with only binary mask $Y\in\{0,1\}^{N}$. Here we utilize the Binary Cross-Entropy (BCE) loss function to compute the segmentation loss, which can be formulated as:
@@ 4.1 Dataset and Experiment Settings
Following previous work [13; 27], we use the mean intersection-over-union (mIoU), and Acc@kIOU as the evaluation metrics. The mIoU is the average of the IoU over all test samples and the Acc@kIOU measures the accuracy of test samples with an IoU higher than the threshold $k$, where $k\in\{0.25,0.5\}$. We use A@25 and A@50 for brevity in some of the following tables.

### v2

@@ Abstract
Referring 3D Segmentation is a visual-language task that segments all points of the specified object from a 3D point cloud described by a sentence of query. Previous works perform a two-stage paradigm, first conducting language-agnostic instance segmentation then matching with given text query. However, the semantic concepts from text query and visual cues are separately interacted during the training, and both instance and semantic labels for each object are required, which is time consuming and human-labor intensive. To mitigate these issues, we propose a novel Referring 3D Segmentation pipeline, Label-Efficient and Single-Stage, dubbed LESS, which is only under the supervision of efficient binary mask. Specifically, we design a Point-Word Cross-Modal Alignment module for aligning the fine-grained features of points and textual embedding. Query Mask Predictor module and Query-Sentence Alignment module are introduced for coarse-grained alignment between masks and query. Furthermore, we propose an area regularization loss, which coarsely reduces irrelevant background predictions on a large scale. Besides, a point-to-point contrastive loss is proposed concentrating on distinguishing points with subtly similar features. Through extensive experiments, we achieve state-of-the-art performance on ScanRefer dataset by surpassing the previous methods about 3.7% mIoU using only binary labels. Code is available at https://github.com/mellody11/LESS.
@@ 1 Introduction
Previous Referring 3D Segmentation [13; 29] mainly leverage a two-stage workaround, as shown in Fig.1 (a). They typically adopt a 3D instance semantic segmentation network to get the instance proposals at the first stage. The predicted instance proposals will be utilized to match with the queries and finally get the final prediction mask according to the matching score. Although those method have achieved remarkable performance, there still exist some problems. First, owing to the large-scale and irregular 3D point clouds, some instance proposals may leave out the target in the pre-segmentation stage. Besides, lacking of linguistic guidance in the segmentation stage fails to focus on the objects that are more essential to the referring task. Moreover, existing Referring 3D Segmentation utilizes both instance labels and semantic labels to segment target proposal rather than binary mask used in referring image segmentation, which is more time consuming and labor intensive.
@@ 2.1 Referring 3D Segmentation
3D visual grounding aims to locate the object within point clouds mentioned by free-form natural language descriptions. Most methods follow a two-stage detection-then-matching pipeline. Initially, they utilize a pre-trained 3D detector [28; 23] or segmenter [15; 33] to extract object representations. Subsequently, these methods align text features with object features to identify the best-matched object. Researchers primarily concentrate on the second stage which involves modeling object relationships and exploring feature fusion between language and objects. Methods employed include multi-modal feature concatenation [3; 1], attention-based multi-modal feature alignment [44; 11], graph neural network-based reasoning [13; 41], and the alignment of visual and language features aided by 2D images [40; 42; 38; 37].Other researchers have investigated single-stage approaches for 3D visual grounding. 3D-SPS [25] views the task as key point selection, progressively identifying keypoints with the guidance of language and directly locates the target. BUTD-DETR [14] employs a transformer decoder [2] to identify described objects using language cues and proposal boxes. Building on this, EDA [36] enhances dense alignment between objects and point clouds by explicitly decoupling textual attributes from sentences.
@@ 2.3 Referring Image Segmentation
Referring image segmentation is a visual task that involves pixel-level segmentation of an image’s target object based on a referring expression. Early approaches [12; 20; 18; 4; 17] utilize CNNs and RNNs to extract features and then perform simple concatenation for multi-modal feature fusion. Subsequent research [16; 19; 9] adopt transformer models for more effective feature extraction and fusion. Recent studies [39; 10; 43] focus on identifying optimal positions for language-vision alignment. Additionally, some methods [34; 5] have enhanced alignment between language and pixel features by designing specialized loss functions. Other approaches [21; 45] treat referring image segmentation as an auto-regressive vector generation problem, creating masks from generated closed vectors.
@@ 3.1 Visual and Text Feature Extractor
The point cloud scene $P\in\mathbb{R}^{N\times 6}$ contains $N$ points in the scene, and each point is represented with six dimensions of RGBXYZ attributes. We first voxelize points into regular voxels and adopt a sparse 3D U-Net [31; 15] to extract point-wise fused feature $F\in\mathbb{R}^{N\times C}$. Here the encoder part of the U-Net has $5$ stages and the feature from the $i$-th encoder stage is denoted as $V_{i}$.
@@
Due to the lack of such rich annotations as [13; 29], it is crucial for our model to learn the relationship between fine-grained word-level features and point-level features in such a point-level segmentation task. Meanwhile, we notice that leveraging the rich convolutional layers in the encoder to excavate multi-modal context is effective way to extract language-aware visual feature. Therefore, we design Point-Word Cross-Modal Alignment (PWCA) module, as shown in the lower left part in Fig.2, which contains a standard cross-attention module [32] and an MLP-Tanh gate. Cross-attention module aligns point-wise and word-wise feature to get the language-aware visual feature. A nonlinear tanh gate is adopted to prevent the fused signal from overwhelming the original signal. Given the multi-scale point cloud features $\{V_{i}\in\mathbb{R}^{N_{i}\times C_{i}}\}_{i=1}^{5}$, we simply project the word feature $W\in\mathbb{R}^{L\times C}$ to $W_{i}\in\mathbb{R}^{L\times C_{i}}$. PWCA can be formulated as follow:
@@ 3.2 Query Mask Predictor
Inspired by [31; 30], Query Mask Predictor (QMP) module, as shown in Fig.2, takes fused feature $F\in\mathbb{R}^{N\times C}$ and learnable queries $Q_{0}\in\mathbb{R}^{K\times C}$ as input and progressively distinguishes the referring target by multi-layer cross-modal transformers. Finally, we extract the proposal masks $M_{m}\in\mathbb{R}^{K\times N}$ based on queries embeddings $Q_{m}\in\mathbb{R}^{K\times C}$ and fused feature $F$.
@@ 3.4 Loss Function
Different from previous work [13; 29], we take the Referring 3D Segmentation task as segmentation task with only binary mask $Y\in\{0,1\}^{N}$. Here we utilize the Binary Cross-Entropy (BCE) loss function to compute the segmentation loss, which can be formulated as:
@@ 4.1 Dataset and Experiment Settings
Following previous work [13; 29], we use the mean intersection-over-union (mIoU), and Acc@kIOU as the evaluation metrics. The mIoU is the average of the IoU over all test samples and the Acc@kIOU measures the accuracy of test samples with an IoU higher than the threshold $k$, where $k\in\{0.25,0.5\}$. We use A@25 and A@50 for brevity in some of the following tables.
@@ 5 Conclusions
6 Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants 62371310, in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515011236, in part by the Stable Support Project of Shenzhen (Project No.20231122122722001).
We also thank the support by the MUR PNRR project FAIR (PE00000013) funded by the NextGenerationEU, the PRIN project CREATIVE (Prot. 2020ZSL9F9), and the EU Horizon project ELIAS (No. 101120237). We also acknowledge the CINECA award under the ISCRA initiative, for the availability of partial HPC resources support.